# Editing the Mind of Giants: An In-Depth Exploration of Pitfalls of Knowledge Editing in Large Language Models

### v1

@@ 3.3 Deterioration of General LLM Abilities
Current evaluation metrics are primarily limited to scenarios where editing is performed only once or infrequently, prompting some studies to extend evaluations to the outcomes after consecutive edits. A study by  (Gupta et al., 2024a) discovers that post-edit models exhibit susceptibility to both gradual forgetting and catastrophic forgetting in sequential editing scenarios. Notably, their findings indicate that the extent of knowledge forgetting is more pronounced in meta-learning-based methods compared to locate-and-edit methods. Additionally, models with parameters modified successively show a decline in performance across various downstream NLP tasks (Gu et al., 2024). Furthermore, perplexity is found to increase after consecutive edits across all parameter-modified methods and different LLMs, and is proposed as another metric to indicate model collapse (Yang et al., 2024). These findings further corroborate that model editing aimed at modifying parameters adversely affects the general capabilities of the original LLMs.
@@
Table 3: Experimental results for portability and generalization. SR: Subject-Replace, RR: Reverse-Relation, OH: One-Hop Accuracy, EW: Edit-wise, IW: Instance-wise, MH: Multi-hop Accuracy, CS: Conflict score, CM: Conflict magnitude. Higher values indicate better performance for all metrics in this table. Results marked with $\dagger$ are obtained in our own experiments, and other results are taken from previous studies.
@@
Table 4: Experimental results for locality. OA: Other-Attribution, DN: Distract-Neighbor, OT: Other-Task, Succ.: Success rate, D: Distortion, IR: Ignore rate, FR: Failure rate, TFD: Tied fact damage. Unless specifically indicated by a downward arrow, higher values signify better performance in those evaluation metrics. Results marked with $\dagger$ are obtained in our own experiments, and other results are taken from previous studies.
@@ 4 Experiments
The experiments were done to evaluate robust generalization and locality (Section  4.1.1 as well as deterioration of general LLM abilities (Section  4.1.2 across different editing methods.
@@ 4.1.1 Robust generalization and locality
We use GPT-J (Wang and Komatsuzaki, 2021) as the baseline model for editing and implement six distinct editing methodologies to assess robust generalization and locality: MEND (meta-learning), ROME and MEMIT (locate-and-edit), SERAC (external memory), and IKE (prompting demonstrations).
Given the overlap in benchmarks for robust generalization and locality, we select a subset for our experiments. Robust generalization is evaluated under single edit and multiple edit settings. Single edit metrics include Subject-Replace, Reverse-Replace, and One-Hop reasoning (Yao et al., 2023a). Multiple edit metrics include multi-hop editing accuracy (Zhong et al., 2023), and Conflict Score and Conflict Magnitude for Reverse Conflict and Composite Conflict respectively (Li et al., 2024b).
For locality, single edit metrics include Other-Attribution, Distract-Neighbor, and Other-Task (Yao et al., 2023b), while multiple edit metrics encompass Success Rate, Distortion, Ignore Rate, and Failure Rate (Li et al., 2024b).
@@ 4.1.2 Deterioration of general LLM abilities
Following the settings of (Gu et al., 2024), we assessed deterioration of general LLM abilities post-editing using six methodologies: ROME, MEMIT, SERAC, MEND, KN, and GRACE. We evaluated general abilities across four NLP downstream tasks: open-domain question answering, sentiment analysis, reasoning, and summarization. These tasks were assessed after 10 to 40 edits on the Zero-Shot Relation Extraction (ZsRE) dataset(Levy et al., 2017), comparing the results against pre-editing benchmarks. More details on the selected downstream tasks are in Appendix B.
@@ 4.2 Experimental Results and Discussion
In general, current editing methodologies exhibit suboptimal performance concerning both robust generalization and locality. Regarding robust generalization, IKE, which leverages prompt demonstrations, demonstrates superior performance over other methodologies in single edit conditions. However, IKE’s performance noticeably declines in multiple edit scenarios, suggesting that prompt demonstrations may become confused when editing multiple logically related facts. Conversely, fine-tuning and meta-learning-based methods are less susceptible to confusion after editing multiple related facts.
Regarding locality, IKE maintains stable performance across metrics in single edit settings. Parameter-modifying methods excel in Other Attribution but decline in other metrics, except MEMIT, which remains stable across all metrics. In multiple edit scenarios, all methods except SERAC show similar performance. In the multiple edit scenario, all methods except SERAC exhibit relatively similar performance. SERAC displays low edit success rate and distortion rate, suggesting its scope classifier does not adopt most edits in this scenario. This may be attributed to its weakness in recovering edited facts, which is crucial in this metric setting.
In terms of general LLM abilities, the number of edits affects methods differently. Meta-learning methods like MEND degrade significantly after 10-20 edits. Locate-and-edit methods such as ROME and KN degrade after 10 edits, while MEMIT remains stable after 40 edits. This disparity can be attributed to MEMIT’s strategy of adjusting parameters across multiple layers, as opposed to ROME’s single-layer edits and KN’s approach of modifying a few neurons. This distribution of parameter modifications across layers helps mitigate deterioration.
@@
(a) Open Domain Question Answering
(b) Sentiment Analysis
(c) Summarization
(d) Reasoning
Figure 4: The experimental results for the deterioration of general abilities were obtained by editing a pretrained LLM using GPT-J with various editing algorithms, including ROME, MEMIT, MEND, KN, SERAC, and GRACE, each applied 10 to 40 times. The edited models were subsequently evaluated on four downstream tasks, including open-domain question answering, sentiment analysis, summarization, and reasoning.
@@ 5.1 Leveraging Information Retrieval and External Memory
Previous research has demonstrated the benefits of utilizing external knowledge bases, as opposed to relying solely on internal knowledge, to guide LLMs in generating content based on a predefined set of facts. These methods effectively separate the factual knowledge stored in LLMs from the inference processes, thus reducing potential biases encoded within the models.
External knowledge bases can include diverse sources such as extensive text corpora, structured tables, or even simple key-value databases. Once these knowledge sources are provided, one can either finetune the LLMs to enhance their ability to retrieve information or employ prompting and in-context learning techniques to query these sources while keeping the model parameters intact. Such approaches not only eliminate the need to verify and edit false factual knowledge within the LLMs but also facilitate the use of attribution and reflection methods. This ensures that the generated content aligns with the predefined external knowledge base, thereby enhancing both accuracy and accountability.
5.2 Improve Understandings of LLMs’ Internal Knowledge Structures
@@
While the identification of factual knowledge storage in LLMs has been extensively explored in recent literature (Meng et al., 2022, 2023; Dai et al., 2022; Hernandez et al., 2024; Geva et al., 2021), the correlation between the location of knowledge and the success rate of model editing remains low (Hase et al., 2023). Additionally, despite evidence suggesting a strong connection between factual knowledge and the feed-forward network layers (Meng et al., 2022; Geva et al., 2021, 2022), recent findings (Li et al., 2024a) indicate that updates to multi-head self-attention layers also lead to improved outcomes. These studies highlight that merely locating fact storage does not fully elucidate the underlying mechanisms of knowledge structures in LLMs. Therefore, further research into how knowledge locations interact with model predictions is essential for advancing the interpretability and controllability of LLMs.
@@
In addition to enhancing the success rate of edits, preserving the general capabilities of LLMs is crucial for assessing the efficacy of model editing methods, as discussed in Section 3.3. Recent breakthroughs in identifying regions within models that correlate with general linguistic abilities (Zhang et al., 2024b) have opened up a direction for future research in model editing. Specifically, by locating these critical areas, it is possible to perform targeted modifications while keeping these regions away from alterations, thereby preventing the deterioration of general abilities. Consequently, advancements in related fields would ensure that edits could be performed without compromising the overall performance of the LLMs, thereby significantly enhancing the specificity and effectiveness of current model editing methods.
5.3 Improve Robustness of Knowledge Editing
Even after successful (achieving fair scores on the existing metrics) edit, the revised model may refuse the modification if the knowledge concerning the altered concept is challenged by extended dialogues. Instead, it might revert to the pre-edit version (reversion), or output an ambiguous answer about the edited concept (confusion). Given the interconnected nature of knowledge, experiments disclose that more popular the knowledge in the benchmark, the easier for the modified model to trace back the original concept (Ma et al., 2024). It highlights the unsatisfying robustness of the existing editing strategies. A more comprehensive understanding how LLMs store and process among different knowledge entities is crucial for a more robust editing. We are also short of specific benchmarks and automated metrics addressing on theses issues. Knowledge-focused editing would not avoid the hallucination inherited from the pre-edit model. TruthX (zhang2024truthx) tries to alleviate hallucination via a parameter-preserved approach by mapping the LLMs internal representation to semantic and truthful spaces and edits the truthfulness in the truthful space. Combination of truthfulness and knowledge adjustment in the same space may provide a practical solution.
@@ 6 Conclusion
Although model editing techniques appear promising for cost-effectively updating knowledge, they still have significant pitfalls. Current editing methods often struggle with making logical inferences based on the edited facts, introducing unintended alterations of non-target knowledge and deterioration in model performance, particularly with parameter-modified methods. Editing techniques that leverage information retrieval can mitigate deviations in model abilities by keeping model parameters intact, as demonstrated in our experiments. Moreover, gaining a deeper understanding of how models store and process knowledge can enhance the controllability of edited facts, leading to greater robustness. We hope our work illuminates potential directions for future improvements in knowledge editing.
@@ A.1 Portability / Generalization
In the single edit scenario, we further classify the methods into two settings:
@@
  • Multi-Hop: This setting assesses the impact of a single edit on more complex, multi-hop reasoning tasks.
For one-hop evaluations, we adopt the methods proposed by (Yao et al., 2023a). These include:
@@
  • Reversed Relation: This metric evaluates the model’s capability to handle reversed relations by filtering for suitable relations such as one-to-one and asking the reverse question to check if the target entity is also updated.
@@
    where $f^{*}$ is the model after editing.
@@
Multiple Edits
In the multiple edits scenario, we test the model’s performance after applying multiple edits. For this, we use the setting and evaluation methods from (Li et al., 2024b). The settings consist of:
@@
    $\left\{\begin{array}[]{l}k_{0}=(s_{1},r_{1},o_{2})\\ k_{n}=(s_{2},r_{1},o_{2})\end{array}\right.$
@@ A.2 Locality
In the single edit scenario for locality, we adopt the methods proposed by (Yao et al., 2023b), including:
  • Other Attribution The modified CounterFact dataset is applied to test whether the non-target attributes of the edited subjects remained the same. For example, if we reset Lionel Messi as a basketball player, his nationality should stay the same.
  • Distract Neighbor Pretend model editing by modifying the neighborhood prompt in the CounterFact. For example, if the original prompt is "Windows 11 is a product of __", the modified prompt would be "Windows 11 is a product of Google. Office 365, developed by __". It testifies whether the model prediction would be "distracted" by the revised prompt.
  • Other Task The edited model is tested on the multiple-choice QA task Physical Interaction QA(PIQA, Bisk et al. (2020)) and the performance is evaluated by accuracy.
@@
We also test the model’s locality in the multiple edits scenario adopting the methods and evaluations from (Li et al., 2024b). The settings consist of:
@@
  • Failure Rate (FR) (Li et al., 2024b):
@@
  • Tied Fact Damage (TDF) (Li et al., 2024b):
@@
  • Neighborhood Score (NS) (Meng et al., 2022): collect a set of "neighbor" subjects and evaluation the success fraction for $\mathbb{P}\left[o^{c}\right]>\mathbb{P}\left[o^{*}\right]$, while the $o^{c}$ denotes the correct facts and $o^{*}$ denotes the false facts.
@@
Appendix B Detailed Explanation of experiments for deterioration of general LLM abilities
We follow the settings of  (Gu et al., 2024) for this part of experiments. Different evaluation metrics were applied for each downstream task: Exact Match for open-domain question answering on the Natural Question dataset  (Kwiatkowski et al., 2019), accuracy for sentiment analysis on the SST2 dataset  (Socher et al., 2013), solve rate for reasoning on the GSM8K dataset  (Cobbe et al., 2021), and ROUGE score for summarization on the SAMSum dataset  (Wang et al., 2019).

### v2

@@ 3.3 Deterioration of General LLM Abilities
Current evaluation metrics are primarily limited to scenarios where editing is performed only once or infrequently, prompting some studies to extend evaluations to the outcomes after consecutive edits. A study by Gupta et al. (2024a) discovers that post-edit models exhibit susceptibility to both gradual forgetting and catastrophic forgetting in sequential editing scenarios. Notably, their findings indicate that the extent of knowledge forgetting is more pronounced in meta-learning-based methods compared to locate-and-edit methods. Additionally, models with parameters modified successively show a decline in performance across various downstream NLP tasks (Gu et al., 2024). Furthermore, perplexity is found to increase after consecutive edits across all parameter-modified methods and different LLMs, and is proposed as another metric to indicate model collapse (Yang et al., 2024). These findings further corroborate that model editing aimed at modifying parameters adversely affects the general capabilities of the original LLMs.
@@
Table 3: Experimental results for portability and generalization. SR: Subject-Replace, RR: Reverse-Relation, OH: One-Hop Accuracy, MH: Multi-hop Accuracy, CS: Conflict score, CM: Conflict magnitude. Higher values indicate better performance for all metrics in this table. Results marked with $\dagger$ are obtained in our own experiments, and other results are taken from previous studies.
@@
Table 4: Experimental results for locality. OA: Other-Attribution, DN: Distract-Neighbor, OT: Other-Task, Succ.: Success rate, D: Distortion, IR: Ignore rate, FR: Failure rate. Unless specifically indicated by a downward arrow, higher values signify better performance in those evaluation metrics. Results marked with $\dagger$ are obtained in our own experiments, and other results are taken from previous studies.
@@ 4 Experiments
The experiments are done to evaluate robust generalization and locality (Section  4.1.1 as well as deterioration of general LLM abilities (Section  4.1.2 across different editing methods.
@@ 4.1 Experimental Setup
Given the variety of benchmarks addressing different challenges in knowledge editing, systematically comparing model performance becomes difficult. To address this, we select the most widely used datasets for each category of pitfalls, ensuring a fair and transparent comparison.
@@ 4.1.1 Robust Generalization and Locality
We use GPT-J (Wang and Komatsuzaki, 2021) as the baseline model for editing and implement six distinct editing methodologies to assess robust generalization and locality: MEND (meta-learning), ROME and MEMIT (locate-and-edit), SERAC (external memory), and IKE (prompting).
Given the overlap in benchmarks for robust generalization and locality, we select a subset for our experiments. The evaluation is divided into two settings: single edit, where only one fact in a reasoning chain is modified, and multiple edits, where several logically connected facts in the chain are altered simultaneously. A detailed description is provided in the Appendix A). Single edit metrics include Subject-Replace, Reverse-Replace, and One-Hop reasoning (Yao et al., 2023). Multiple edit metrics include multi-hop editing accuracy (Zhong et al., 2023), and Conflict Score and Conflict Magnitude for Reverse Conflict and Composite Conflict respectively (Li et al., 2024). For locality, single edit metrics include Other-Attribution, Distract-Neighbor, and Other-Task (Yao et al., 2023), while multiple edit metrics encompass Success Rate, Distortion, Ignore Rate, and Failure Rate (Li et al., 2024).
(a) Open Domain Question Answering
(b) Summarization
(c) Sentiment Analysis
(d) Reasoning
Figure 4: The experimental results for the deterioration of general abilities were obtained by editing GPT-J with various editing algorithms, including ROME, MEMIT, MEND, KN, SERAC, and GRACE, each applied 10 to 40 times. The edited models were subsequently evaluated on four downstream tasks, including open-domain question answering, sentiment analysis, summarization, and reasoning. The results for SERAC and GRACE are overlapping.
@@ 4.1.2 Deterioration of General LLM Abilities
Following the settings of (Gu et al., 2024), we assess deterioration of general LLM abilities post-editing using six methodologies: ROME, MEMIT, SERAC, MEND, KN, and GRACE. We evaluate general abilities across four NLP downstream tasks: open-domain question answering, sentiment analysis, reasoning, and summarization. These tasks are assessed after 10 to 40 edits on the Zero-Shot Relation Extraction (ZsRE) dataset(Levy et al., 2017), comparing the results against pre-editing benchmarks. More details on the selected downstream tasks are in Appendix B.
@@ 4.2 Experimental Results and Discussion
In general, current editing methodologies show suboptimal performance in both robust generalization and locality. Regarding robust generalization (Table 3), IKE, which leverages prompt demonstrations, excels in single edit but declines with multiple edits. This suggests that prompt demonstrations may become confused when editing multiple logically related facts. Conversely, fine-tuning and meta-learning-based methods are less susceptible to confusion after editing multiple related facts.
Regarding locality (Table 4), IKE maintains stable performance across metrics in single edit settings. Parameter-modifying methods excel in Other Attribution but decline in other metrics, except MEMIT, which remains stable across all metrics. In multiple edit scenarios, all methods except SERAC show similar performance. In the multiple edit scenario, all methods except SERAC exhibit relatively similar performance. SERAC displays low edit success rate and distortion rate, suggesting its scope classifier does not adopt most edits in this scenario. This may be attributed to its weakness in recovering edited facts, which is crucial in this metric setting.
In terms of general LLM abilities (Figure 4), the number of edits affects methods differently. Meta-learning methods like MEND degrade significantly after 10-20 edits. Locate-and-edit methods such as ROME and KN degrade after 10 edits, while MEMIT remains stable after 40 edits. This disparity can be attributed to MEMIT’s strategy of adjusting parameters across multiple layers, as opposed to ROME’s single-layer edits and KN’s approach of modifying a few neurons. This distribution of parameter modifications across layers may help mitigate deterioration.
@@ 5.1 Leveraging Information Retrieval and External Memory
Research shows that using external knowledge bases, rather than relying solely on internal knowledge, benefits LLMs by guiding content generation based on predefined facts. External knowledge sources, such as text corpora, structured tables, or key-value databases, can be utilized either to finetune LLMs for improved information retrieval or to employ prompting techniques for querying these sources. These approaches separate factual knowledge from inference process, thus preserves the original model parameters and minimizes post-editing damage. Moreover, they ensure that generated content aligns with predefined knowledge bases, thereby enhancing accountability and accuracy.
5.2 Improving Understandings of LLMs’ Internal Knowledge Structures
@@
While identifying where factual knowledge is stored in LLMs has been extensively explored (Meng et al., 2022, 2023; Dai et al., 2022; Hernandez et al., 2024; Geva et al., 2021), the correlation between knowledge location and editing success remains low (Hase et al., 2023). Additionally, despite evidence suggesting a strong connection between factual knowledge and the feed-forward network layers (Meng et al., 2022; Geva et al., 2021, 2022), recent findings indicate that updates to multi-head self-attention layers also improve outcomes (Li et al., 2023). This suggests that locating fact storage alone doesn’t fully explain knowledge structures in LLMs. Further research is needed to understand how knowledge locations interact with model predictions in order to enhance LLM interpretability and controllability.
@@
Preserving LLMs’ general capabilities is also crucial for model editing, as discussed in Section 3.3. Recent breakthroughs in identifying regions within models that correlate with general linguistic abilities have opened up a direction for future research in model editing (Zhang et al., 2024b). By making targeted modifications, we can potentially prevent the deterioration of general abilities and improve the specificity and effectiveness of model editing methods.
5.3 Improving Robustness of Knowledge Editing
Even after achieving fair scores on the existing metrics, models may revert to pre-edit versions or provide ambiguous answers if the altered knowledge is conflicted with inherited concepts. Experiments show that more popular knowledge is easier for modified models to revert to (Ma et al., 2024), indicating the lack of robustness in current editing strategies. A deeper understanding of how LLMs store and process interconnected knowledge entities is crucial for more robust editing and warrants future research.
@@ 6 Conclusion
Although model editing techniques appear promising for cost-effectively updating knowledge, they still have significant pitfalls. Current editing methods often struggle with making logical inferences based on the edited facts, introducing unintended alterations of non-target knowledge and deterioration in model performance, particularly with parameter-modified methods. By harnessing information retrieval techniques and delving into how models store and process knowledge, deviations in model abilities can be mitigated, and the controllability of edited facts can be enhanced, ultimately leading to greater robustness. We hope our work illuminates potential directions for future improvements in knowledge editing.
Limitations
The field of knowledge editing is advancing at an impressive pace, with numerous innovations in editing methodologies and evaluation metrics being proposed. Despite our efforts to collect and organize previous work, some contributions may not be included in this paper. However, we will continue to monitor the latest developments in this field and update our GitHub repository with recent related works.
Acknowledgments
We thank the reviewers for their insightful comments. This work was financially supported by the National Science and Technology Council (NSTC) in Taiwan, under Grants 111-$2$-E-002-013-MY3 and 112-2223-E002-012-MY5. We thank to National Center for High-performance Computing (NCHC) of National Applied Research Laboratories (NARLabs) in Taiwan for providing computational and storage resources.
@@ A.1 Portability / Generalization
In the single edit scenario, we modify only one fact in the logical chain with each edit. Let $Z_{e}=\{(x_{e},y_{e})\ |\ f_{\theta_{0}}(x_{e})\neq y_{e}\}$ be the set where only a single fact is edited in each logical chain. Single edit is conducted as:
$f_{\theta_{e}}(x_{e})=y_{e},\forall(x_{e},y_{e})\in Z_{e}$ (5)
This part consists of:
@@
For one-hop evaluations, we adopt the methods proposed by (Yao et al., 2023). These include:
@@
  • Reversed Relation: This metric evaluates the model’s capability to handle reversed relations by filtering for suitable relations (e.g., one-to-one relation) and asking the reverse question to check if the target entity is also updated.
@@
Multiple Edits
In the multiple edits scenario, we evaluate the model’s performance after applying several logically related edits. Let $Z_{e}=\{(x_{ei},y_{ei})\ |\ f_{\theta_{0}}(x_{ei})\neq y_{ei}\}$ represent a set of logically related facts within a reasoning chain intended to be modified. Multiple edits are performed by altering several facts within this chain:
$f_{\theta_{e}}(x_{ei})=y_{ei},\forall(x_{ei},y_{ei})\in Z_{e}$ (6)
This part consists of:
  • Multi-Hop editing: Evaluate whether the model can infer edited knowledge in multi-hop questions.
  • Conflict editing: Assess how the model handles multiple conflicting edits.
@@
    where $f^{*}$ is the model after editing, $s$ refers to the edited subject, and $o$ refers to target object.
@@
As for Conflict editing, we use the setting and evaluation methods from (Li et al., 2024). The settings consist of:
@@
    $\left\{\begin{array}[]{l}k_{o}=(s_{1},r_{1},o_{2})\\ k_{n}=(s_{2},r_{1},o_{2})\end{array}\right.$
    where $k_{o}$ refers to old knowledge, and $k_{n}$ refers to new knowledge.
@@ A.2 Locality
In the single edit scenario for locality, we adopt the methods proposed by (Yao et al., 2023), including:
  • Other Attribution (OA): The modified ZsRE and CounterFact datasets are applied to test whether the non-target attributes of the edited subjects remained the same. For example, if we reset Lionel Messi as a basketball player, his nationality should stay the same.
  • Distract Neighbor (DN): Previous studies indicate that if edit cases are concatenated with unrelated context, the model tends to output content related to the edit cases. For example, if the original prompt is "Windows 11 is a product of __", an edit case is added in front and be "Windows 11 is a product of Google. Office 365, developed by __". It testifies whether the model prediction would be "distracted" by the edit case.
  • Other Task (OT) The edited model is tested on the multiple-choice QA task Physical Interaction QA (PIQA, Bisk et al. (2020)) and the performance is evaluated by accuracy.
@@
We also test the model’s locality in the multiple edits scenario by adopting the methods and evaluations from (Li et al., 2024). The settings consist of:
@@
    measures the extent to which objects in Obj set (excluding the target object $o_{1}$) are disregarded or overlooked after the process of knowledge editing.
  • Failure Rate (FR) (Li et al., 2024):
@@
    calculates the rate when Ignore Rate > 0.5
  • Tied Fact Damage (TDF) (Li et al., 2024):
@@
  • Neighborhood Score (NS) (Meng et al., 2022): collect a set of "neighborhood" subjects and evaluate the success fraction for $\mathbb{P}\left[o^{c}\right]>\mathbb{P}\left[o^{*}\right]$, while the $o^{c}$ denotes the correct facts and $o^{*}$ denotes the false facts.
@@
Appendix B Detailed Experimental Details of the Deterioration of General LLM Abilities
We follow the settings of  (Gu et al., 2024) for this part of experiments. Different evaluation metrics were applied for each downstream task: Exact Match for open-domain question answering on the Natural Question dataset  (Kwiatkowski et al., 2019), accuracy for sentiment analysis on the SST2 dataset  (Socher et al., 2013), solve rate for reasoning on the GSM8K dataset  (Cobbe et al., 2021), and ROUGE score for summarization on the SAMSum dataset  (Gliwa et al., 2019).